# Two Views Are Better than One: Monocular 3D Pose Estimation with Multiview Consistency

### v2

@@ 4.4 Fine-tuning without available 3D data
Again we add the consistency loss as a regularization term resulting in the total loss in Equation 8. Through experimentation, we found that when two camera views are available, we achieve the best performance, with $\lambda_{2D_{\text{reproj}}}=1$ and $\lambda_{\text{con}}=0.3$,
@@
By fine-tuning the MotionBERT [46] model with the losses in Equations 7 and 8, using two camera views as described in Section 4.3, we achieve the results in Table 1.
The results in Tab. 1 highlight the impact of the consistency loss regularization term, particularly in scenarios where ground truth 3D information is absent. This regularization term leads to a substantial improvement in MPJPE, demonstrating a reduction of $39.2$mm compared to relying solely on the reprojection loss. Visualizing 3D predictions from models with and without the consistency loss, Figure 3(a), we see the same substantial increase in accuracy when using the consistency loss.
@@
[…]
(a) Visual comparison of predictions in green and the ground truth pose in blue. The magnitude of errors, measured in millimeters and indicated at the top, highlights the superiority of our consistency loss $\mathcal{L}_{\text{2D}_{\text{con}}}$ in achieving more accurate results. The notable improvement is especially evident in the bottom row, where the method employing our consistency loss successfully captures the complex movement.
Right $+$
view $x$
@@
(b) MPJPE and PA-MPJPE for different combinations of two views. The view “right” is included in all combinations. All experiments have been conducted with $\mathcal{L}_{\text{2D}_{\text{con}}}$, $\lambda_{\text{con}}=1$. It is clear that the two-view combination matters with views 1 + right and view 2 + right achieving substantially lower errors.
Figure 3: 3(a): Visualization of the effect of our loss. 3(b): Overview of which views to use.
We believe this improvement is this big because the consistency loss has improved the network’s ability to resolve ambiguities during the process of lifting 2D to 3D from a single view. In addition, it proves beneficial in situations where joints might be occluded in one of the views, enhancing the overall robustness of the model.
However, a closer examination of the Procrustes aligned joint error in Table 1 reveals an interesting observation. Fine-tuning the model solely on 2D body keypoints results in an increase in error. This phenomenon could be attributed to the inherent ambiguity associated with multiple different 3D poses that can reproject to the same 2D body pose. Consequently, the model may struggle to provide accurate depth estimates of joint locations, as highlighted in the work by Ingwersen et al. [11]. This underscores the importance of the consistency loss in mitigating such challenges and emphasizes its role in refining the model’s performance in the absence of ground truth 3D data.
@@
Table 2 shows results from our method with consistency loss compared with other semi-supervised methods trained on the Human3.6M [13] dataset. For the evaluation we adopt the same protocol as Iqbal et al. [14] using 3D supervision from subject (S1) and use other the subjects (S5, S6, S7, S8) for weak supervision with our consistency loss, $\mathcal{L}_{\text{2D}_{\text{con}}}$ from Eq. 8. For evaluation, we use the standard protocol testing on subjects S9 and S11 [13, 14].
Table 2: Comparison with state-of-the-art Semi-Supervised methods on the Human3.6M[13] dataset using only 3D data from Subject 1 for training.
@@
Rodhin et al. (ECCV’18) [35] $131.7$ $98.2$
Pavlakos et al. (ICCV’19) [32] $110.7$ $74.5$
Li et al. (ICCV’19) [22] $88.8$ $66.5$
Rodhin et al. (CVPR’18) [36] - $65.1$
Kocabas et al. (CVPR’19) [20] - $60.2$
Iqbal et al. (CVPR’20) [14] $62.8$ $51.4$
Roy et al. (3DV’22) [37] $60.8$ $48.4$
Ours $58.9$ $43.6$
The data depicted in Table 2 demonstrates that our novel consistency loss significantly improves performance in scenarios with limited 3D data with unlabeled multiview data. A key innovation of our approach, compared to Iqbal et al. [14], lies in incorporating the temporal aspect by computing a similarity transform per sequence instead of per frame, and not requiring known camera intrinsics. This advancement establishes a new state-of-the-art in semi-supervised performance on the Human3.6M dataset [13].
@@ 4.6 How many views do we need?
Examining the experiments carried out in Sections 4.3 and 4.4, a logical inquiry arises regarding the scalability of the results when more than two views are incorporated into the experiments. To investigate the correlation between the number of views and performance, we have calculated the results for scenarios where one to seven views are available, encompassing the total number of views in the SportsPose dataset [12].
@@ 4.7 More views or more data?
In the experiment analogous to 2D supervision illustrated in Figure 4(a), an examination of the results without the consistency loss in Figure 4(c) reveals that neither MPJPE nor PA-MPJPE exhibit improvement with the addition of more training data through an increased number of views. The observed plateau after two views contradicts the substantial accuracy increases depicted in Figure 4(a), suggesting that these improvements are attributed to the introduction of our consistency loss.
@@
0.50.50.50.511111.51.51.51.522222.52.52.52.533333.53.53.53.544444.54.54.54.555555.55.55.55.566666.56.56.56.577777.57.57.57.50055551010101015151515202020202525252530303030ViewsError (mm)MPJPEPA-MPJPE
@@ 4.8 Which views to choose
As the ambiguities of a 3D pose can theoretically be solved when observing from any two different views, a question arises if there is a practical advantage of certain viewpoints. This was investigated in Figure 3(b) where the model was trained using the “right” view in combination with all other available views, where View 1 corresponds to the perspective positioned 90 degrees relative to the view facing the back of the subject.
Analyzing the errors depicted in Figure 3(b), it is evident that the choice of the view for multiview supervision significantly influences the outcomes of using our consistency loss. This aligns with intuition, as certain views are more effective in resolving ambiguities and identifying occluded joints, while others may not contribute new information. The results indicate that an optimal configuration involves using views from two sides that are 90 degrees apart when only two cameras are available.
@@ 5 Discussion and conclusion
Limitations. While our results underscore a notable improvement in accuracy achieved through the implementation of our consistency loss, it is important to acknowledge the limitations of our method. As Figure 3(b) demonstrates, the effect of the consistency loss is dependent on which views are used during training, with the least favorable combination resulting in performance comparable to using a single camera view. However, a substantial increase in accuracy is evident in four out of five combinations.
Furthermore, it is essential to highlight that an excessive emphasis on the consistency loss, indicated by a large $\lambda_{\text{con}}$ value, can lead to a degenerate solution. Specifically, an optimal solution to Equation 2 may manifest as predicting an identical position of all joints, emphasizing the need for careful consideration when setting this parameter.
@@
It is worth noting that incorporating the proposed consistency loss necessitates temporal synchronization of pose sequences from different views. Desynchronisation will appear as correlated noise in the 2D predictions that in the worst case can confuse the model, especially for fast movements. This requirement imposes constraints on the camera system used for data capture. Using two cameras capturing frames with logged timestamps, it is possible to manually identify the same point in time in both sequences, or to use the audio to time-synchronize the views after acquisition [23]. However, if using Android smartphones as cameras the frame-synchronization can be obtained using an app and WiFi hotspots as described by Akhmetyanov et al. [1].
Conclusion. We present a novel method to enhance monocular 3D human pose estimation performance. By incorporating our multiview consistency loss during training in scenarios where 3D data is unavailable, we achieve notable performance improvements when compared to relying solely on a 2D reprojection loss or no fine-tuning without requiring knowledge of the camera’s extrinsic or intrinsic parameters.
We demonstrate the efficacy of the proposed consistency loss by evaluating it on the Human3.6M [13] and SportsPose [12] dataset. Following the semi-supervised protocol for Human3.6M, we advance the precision and achieve state-of-the-art precision.

### v3

@@ 4.3 Fine-tuning without 3D data
We also add our consistency loss as regularization
@@
Through experimentation, we found that we achieve the best performance, with $\lambda_{2D_{\text{reproj}}}=1$ and $\lambda_{\text{con}}=0.3$, and use these values unless otherwise specified.
Using the same two views as described in Section 4.2 we fine-tune on the SportsPose dataset with the loss from Equation 7 or Equation 8 and present the results in Table 2. Here we also present results with a variant of our consistency loss $\mathcal{L}_{\text{2D}_{\text{con}}^{\text{frame}}}$, where the optimal similarity transform is computed per frame, to allow for moving cameras. In Tab. 2 we present results on fine-tuning on the SkiPose dataset only with $\mathcal{L}_{\text{2D}_{\text{con}}^{\text{frame}}}$, as the cameras are not fixed.
Tables 2 and 2 highlights the impact of our consistency loss, particularly when 3D supervision is unavailable. The addition of the consistency loss on SportsPose leads to a substantial improvement in MPJPE, demonstrating a reduction of $39.2$mm compared to relying solely on the reprojection loss. Figure 3 shows predictions from models with and without the consistency loss and we see the same substantial increase in accuracy when using the consistency loss.
@@
RGB MotionBERT [55] $\mathcal{L}_{\text{2D}}$ $\mathcal{L}_{\text{2D}_{\text{con}}}$, ours
Figure 3: Visual comparison of predictions in green and the ground truth pose in blue. The magnitude of errors, measured in millimeters and indicated at the top, highlights the superiority of our consistency loss $\mathcal{L}_{\text{2D}_{\text{con}}}$ in achieving more accurate results. The notable improvement is especially evident in the bottom row, where the method employing our consistency loss successfully captures the complex movement.
Table 3: MPJPE and PA-MPJPE for different combinations of two views. The view “right” is included in all combinations. All experiments have been conducted with the loss $\mathcal{L}_{\text{2D}_{\text{con}}}$ and $\lambda_{\text{con}}=1$. It is clear that the two-view combination matters with right + view 1 and right + view 2 achieve substantially lower errors.
Right $+$ view $x$
@@
The improvement for 2D data is this big because the consistency loss has improved the network’s ability to resolve ambiguities during the process of lifting 2D to 3D from a single view. In addition, our method proves beneficial in situations where joints might be occluded in one of the views, enhancing the overall robustness of the model.
However, a closer examination of the PA-MPJPE in Table 2 reveals an interesting observation. Fine-tuning the model solely on 2D body keypoints based on Eq. 7 results in an increase in error. A likely cause is that the 3D points that reproject to the same 2D points are not unique. Consequently, the model may struggle to provide accurate depth estimates of joint locations, as highlighted by Ingwersen et al. [11]. This underscores the importance of the consistency loss in mitigating such challenges and emphasizes its role in refining the performance of the model without ground truth 3D data.
@@
Table 4 shows results from our method with consistency loss compared with other semi-supervised methods trained on the Human3.6M [13] dataset. For the evaluation, we adopt the same protocol as Iqbal et al. [14] using 3D supervision from S1 and use S5, S6, S7, S8 with our consistency loss, $\mathcal{L}_{\text{2D}_{\text{con}}}$, from Eq. 8. For evaluation, we use the standard protocol testing on subjects S9 and S11 [13, 14].
Table 4: Comparison with reported state-of-the-art Semi-Supervised methods trained from scratch on the Human3.6M [13] dataset using only 3D data from S1 and weak supervision on S5, S6, S7, S8 during training. Unreported values are marked with "-". All 2D poses are predicted.
@@
Rhodin et al. [39] (ECCV’18) $131.7$ $98.2$
Pavlakos et al. [34] (ICCV’19) $110.7$ $74.5$
Li et al. [26] (ICCV’19) $88.8$ $66.5$
Rhodin et al. [40] (CVPR’18) - $65.1$
Kocabas et al. [23] (CVPR’19) - $60.2$
Iqbal et al. [14] (CVPR’20) $62.8$ $51.4$
Roy et al. [41] (3DV’22) $60.8$ $48.4$
Li and Pun [25] (AAAI’23) $51.9$
Ours, $\mathcal{L}_{\text{2D}_{\text{con}}^{\text{frame}}}$ $52.1$ $41.0$
Ours, $\mathcal{L}_{\text{2D}_{\text{con}}}$ 8 $50.5$ $40.4$
Table 4 demonstrates that our novel consistency loss significantly improves performance in scenarios where 3D data is limited and with unlabeled multiview data. A key innovation of our approach, compared to Iqbal et al. [14], lies in incorporating the temporal aspect by computing a similarity transform per sequence instead of per frame, and not requiring known camera intrinsics. This advancement establishes a new state-of-the-art in semi-supervised performance on the Human3.6M dataset [13]. Recent work [37, 7] obtain very similar performance under the same weakly supervised protocol, but they utilize ground truth 2D poses, making the results not directly comparable.
@@ 4.5 How many views do we need?
Examining the experiments carried out in Sections 4.2 and 4.3, a logical inquiry arises regarding the scalability of the results when more than two views are incorporated into the experiments. To investigate the correlation between the number of views and performance, we have calculated the results for scenarios where one to seven views are available, encompassing the total number of views in the SportsPose dataset [12].
@@ 4.6 More views or more data?
Figure 4(c) shows the experiment analogous to 2D supervision illustrated in Figure 4(a), but without the consistency loss. It reveals that neither MPJPE nor PA-MPJPE exhibit improvement with the addition of more training data when incrasing the number of views. The observed plateau after two views contradicts the substantial accuracy increases depicted in Figure 4(a), suggesting that these improvements are attributed to the introduction of our consistency loss.
@@
22224444666600101010102020202030303030Number of viewsError (mm)MPJPEPA-MPJPE
@@ 4.7 Which views to choose
As the ambiguities of a 3D pose can theoretically be solved when observing from any two different views, a question arises if there is a practical advantage of certain viewpoints. This was investigated in Table 3 where the model was trained using the “right” view in combination with all other available views, where View 1 corresponds to the perspective positioned 90 degrees relative to the view facing the back of the subject.
Analyzing the errors depicted in Table 3, it is evident that the choice of the view for multiview supervision significantly influences the outcomes of using our consistency loss. This aligns with intuition, as certain views are more effective in resolving ambiguities and identifying occluded joints, while others may not contribute new information. The results indicate that an optimal configuration involves using views from two sides that are 90 degrees apart when only two cameras are available.
@@ 5 Discussion and conclusion
Limitations. While our results underscore a notable improvement in accuracy achieved through the implementation of our consistency loss, it is important to use the method with care. As Table 3 demonstrates, the effect of the consistency loss depends on which views are used during training, with the least favorable combination resulting in performance comparable to using a single camera view. However, accuracy increases substantially in four out of five combinations.
Furthermore, emphasizing the consistency loss too much, indicated by a large $\lambda_{\text{con}}$ value, can lead to a degenerate solution. Specifically, an optimal solution to Equation 2 may predict the same position of all joints.
@@
It is worth noting that incorporating the proposed consistency loss necessitates temporal synchronization of pose sequences from different views. Desynchronisation will appear as correlated noise in the 2D predictions that in the worst case can confuse the model, especially for fast movements. This requirement imposes constraints on the camera system used for data capture. Using two cameras capturing frames with logged timestamps, it is possible to manually identify the same point in time in both sequences, or to use the audio to time-synchronize the views after acquisition [27]. However, if using Android smartphones as cameras, the frame-synchronization can be obtained using an app and WiFi hotspots as described by Akhmetyanov et al. [1].
Conclusion. We present a novel method to enhance monocular 3D human pose estimation performance. By incorporating our multiview consistency loss during training in scenarios where 3D data is unavailable, we achieve notable performance improvements when compared to relying solely on 2D reprojection loss or no fine-tuning without requiring knowledge of the camera’s extrinsic or intrinsic parameters.
We demonstrate the efficacy of the proposed consistency loss by evaluating it on the Human3.6M [13], SkiPose [44], and SportsPose [12] datasets. Following the semi-supervised protocol for Human3.6M, we advance the weakly supervised state-of-the-art precision.